OpenStudy (anonymous):

Use retriceociative property of multiplycation to place parentheses in two different ways?

OpenStudy (anonymous):

mnq

OpenStudy (anonymous):

(x*y)*z = x*(y*z) its exactly the same as addition

OpenStudy (anonymous):

mnq = m(nq) = (mn)q

OpenStudy (anonymous):

thanks polpak (:

OpenStudy (anonymous):

what about if there's a numer in ti like 4st ): ?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

would it be 4(st) and (4s)t

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

ok do you understand distributive property-multiplycation?